Find relation between x, y and z in 2/log x=3/log y=4/log z

Grade:12th pass

1 Answers

Aditya Gupta
4 years ago
hello rebel.
let 2/log x=3/log y=4/log z= 1/k
so, x= e^2k, y= e^3k and z= e^4k
so, y^2= e^6k= e^2k*e^4k= xz
or y^2= xz or y/x= z/y.
so x, y, z are in GP.
kindly approve :)